### v2

# An Overview of Some Single Machine Scheduling Problems: Polynomial Algorithms, Complexity and Approximability

Abstract:
Since the publication of the first scheduling paper in 1954, a huge number of works dealing with different types of single machine problems appeared. They addressed many heuristics and enumerative procedures, complexity results or structural properties of certain problems. Regarding surveys, often particular subjects like special objective functions are discussed, or more general scheduling problems were surveyed, where a substantial part is devoted to single machine problems. In this paper we present some results on polynomial algorithms, complexity and approximation issues, where the main focus is on results, which have been published during the last decades in papers, where at least one of the first two authors of this paper was involved. We hope that the reviewed results will stimulate further investigation in related research fields.
Keywords: Single machine scheduling; Polynomial algorithms; Complexity; Approximation algorithms
MSC classification: 90B35, 90C59
1 Introduction
Scheduling problems arise in many real-life applications including our everyday life. We commonly desire to optimize our time arranging different activities, where the time is limited and also the resources that we use are scarce. We wish to optimize some of our personal criteria but we also depend on the availability of the latter resources. As a simple but very common daily example, consider an apartment with a single bathroom, where a small family lives. The family members gets up at some fixed time, and all of them need to spend some time in the bathroom, which is a scarce resource (or machine). Typically such a (renewable) resource may serve a single object or job, a family member in our example. Every member of the family needs to leave the apartment at some fixed time to reach his/her job or school without delay. This time can be different for different members of the family. Here arises a small optimization problem that asks to schedule the time interval for each member of the family so as to meet the deadline of each member of the family.
Generally, in a scheduling problem we are given jobs or tasks which are some requests that need to be performed by some resource, a machine or a processor. In our small example, bathroom requirements for each member of the family are jobs and the bathroom is a machine. We can give a vast amount of examples of jobs and machines in our practical life; e.g., a job in a factory or a software application in a computer or on a smart phone, or a lesson at school. Here machines, computers, smart phones, classrooms and teachers are resource examples of. Every jobs requires some time units on a resource, whereas a resource is able to process at most one job at a time. Indeed, one classroom cannot be used for two different classes simultaneously or a a teacher cannot give two or more classes at the same time. There is a restriction on the whole time window during which a factory or a teacher can work. It is a scheduling task to arrange the order for processing the jobs. Often there is some minimization or maximization criterion called an objective function. For example, a very common criterion is to minimize the maximum job or machine completion time or the makespan which is commonly denoted by $C_{\max}$.
We shall denote the given set of $n$ jobs by $\{1,2,\dots,n\}$. The processing time of job $j$ is denoted by $p_{j}$. Job $j$ may have some additional characteristics. For example, the time moment from when job $j$ is available to be processed is called its release or readiness time and is commonly denoted by $r_{j}$. In many applications, a job, once completed by the machine, needs to be delivered to the customer by an independent transportation unit, which may not be a scarce resource in a given model. Such a delivery time (or tail) of job $j$ is commonly denoted by $q_{j}$. The job delivery takes no machine time (a job is delivered by an independent agent to a customer), though $q_{j}$ contributes to the completion time of job $j$. Hence, the makespan crucially depends on the job delivery times. The due date $d_{j}$ is the desirable completion time of job $j$. Both above parameters specify the urgency of job $j$: the more is $q_{j}$, the more urgent is job $j$; likewise, the less is $d_{j}$, the more urgent is job $j$.
All above parameters are normally non-negative integral numbers. In some cases, job $j$ is not allowed to be completed after its due date. Then the term deadline instead of “due date” is commonly used. Thus, if deadlines are given, our objective is to find a schedule in which all jobs are completed before or at their deadlines. With due dates, our objective will be to minimize or maximize some due date oriented objective function. Let the lateness of a job in a schedule be the difference between its completion time and its due date. One of the common objective functions is the maximal job lateness $L_{\max}$ (which we have to minimize). Another common objective function is the number of jobs completed at or before their due dates $\sum U_{j}$ which we have to maximize ($U_{j}$ is a 0-1 variable indicating whether job $j$ is completed on time or not in the schedule). We may allow to omit some jobs in the case when the total workload is more than the machine capacity. In many applications, job due dates are important since, if a job is not completed by its due date then its benefit or functionality becomes neglectable. Hence, we may wish to maximize the throughput, i.e.,the number of jobs completed on time by their due dates. We may also consider more general criteria to maximize some profits. The profit of job $j$ can be expressed by its weight which is commonly denoted by $w_{j}$. Job preemptions might be allowed or not. In a preemptive scheduling problem, a job can be interrupted on the machine and later resumed on the machine. In a non-preemptive scheduling problem, a job $J_{j}$, once scheduled on the machine, should be processed without any interruption until its full completion on the machine. The machine is called idle at some moment $t$ if no job is processed by the machine at that moment, otherwise it is said to be busy.
Problem restrictions on the way in which the jobs are to be performed by machines define the set of feasible solutions. For example, a machine can process at most one job simultaneously; the precedence relations defined on the set of jobs, restrict the relative order in which the jobs can be scheduled. Scheduling problems are sometimes referred to as sequencing problems since the jobs need to be scheduled in a sequential manner (as no job overlapping is permitted). The precedence relations between the jobs can be represented by directed graphs which are often referred to as precedence (task) graphs. In such graphs, nodes represent jobs, and there is an arc $(i,j)$ if and only if job $i$ immediately precedes job $j$.
A schedule prescribes to each job the starting time on the machine creating in this way a total order on the set of jobs by sequencing the jobs on the machine. Denote by $s^{S}_{j}$ ($c^{S}_{j}$, respectively) the starting time (completion time, respectively) of job $j$ in a schedule $S$. A feasible schedule is a schedule which satisfies all restrictions of the given problem. Since all problems imply resource restrictions, in any feasible schedule, if job $j$ is scheduled on the machine, then $s_{j}^{S}\geq s_{i}^{S}+p_{i}$, for any job $i$ scheduled earlier on the machine in $S$ (recal that the machine can handle only one job at any time). Besides, if $j$ has a release time then it cannot start before this time, i.e., $s_{j}^{S}\geq r_{j}$; note that $c^{S}_{j}=s^{S}_{j}+p_{j}$ whenever no job preemption is allowed. An optimal schedule is a one which minimizes the given objective (goal) function.
Graham’s three-field notation is a common way to abbreviate a scheduling problem. The first, the second and the third fields define the machine environment, the job characteristics and the objective function, respectively. In a single-machine (single-processor) scheduling problem, a “1” is used in the first field in the three-field notation to specify a single machine environment. For example, $1|r_{j},q_{j}|C_{\max}$ represents the single machine scheduling problem with release times and tails to minimize the maximum full job completion time (the makespan).
There exist a huge number of papers dealing with single machine problems. In particular, many heuristics and metaheuristic as well as enumerative algorithms have been suggested and tested for special single machine problems. Regarding surveys, the authors often restrict to special subjects and/or consider more general problem settings. As an example, there exist three very detailed surveys on scheduling problems with setup times or costs, see [4, 5, 6]. For problems with so-called $p$-batching (i.e., several jobs can be processed in parallel in a batch), a survey has been given by Fowler and Mönch [18]. The case of equal processing times has been surveyed by Kravchenko and Werner [32].
Regarding single machine problems, often surveys are given for particular objective functions. As a few examples, we can mention e.g. a survey on the total weighted tardiness problem [1], on the weighted number of tardy jobs problem [2] or again for more general problems, a survey on scheduling with late work criteria [46], or more recently on early and late work scheduling [47], where a substantial part is dedicated to single machine scheduling. While mostly a minimization problem is considered, Gafarov et al. [21] reviewed algorithms and complexity issues for single machine maximization problems.
Finally we note that single machine problems are also important for some practical multi-machine scheduling applications. Often there exists a bottleneck machine in a manufacturing environment. If one can solve the scheduling problem on the bottleneck machine optimally, this is at least a very good base for the overall schedule.
In this paper, we review some results related to polynomial algorithms, the complexity and approximability issues. We explicitly mention that we do not survey metaheuristic and enumerative algorithms. In addition to some necessary basic results, the main focus is on the results which have been presented in the past in papers, where the authors of this work participated.
The rest of this paper is organized as follows. In Section 2, we review some necessary basics for single machine scheduling problems. Then, in Section 3, we discuss polynomial algorithms, complexity and approximability issues for different classes of single machine scheduling problems. We also outline some subjects for future research, partially also for more general multi-machine scheduling problems. The paper ends with some concluding remarks in Section 4.
2 Some Basics
In this section, we give some basic notations, concepts and properties that have been used in the study of single machine scheduling problems. Single machine scheduling problems are trivial if every job has a single parameter, its processing time. For minimizing the makespan, it suffices to construct any feasible schedule without a machine idle time or a gap including jobs in any order:
A gap on a machine is a longest time interval during which this machine is idle, i.e., it processes no job. It can be assumed that there occurs a 0-length gap $(c_{j},t_{i})$ if job $i$ starts at its release time at the the completion time of job $j$.
A block is a consecutive part in a schedule which is preceded and succeeded by a gap.
Single machine scheduling problems get less trivial if jobs are not simultaneously released, i.e., they have also different release times. Then, to avoid the creation of any redundant gap, we just need to keep the machine busy whenever this is possible. It is easy to see that this goal will be achieved if we just order the jobs according to non-decreasing release times. After this preprocessing step, the jobs can be included in this order, again, leaving no avoidable gap. This procedure will take $O(n\log n)$ time due to the preprocessing step.
If both, job release and due dates are given, then it makes sense to minimize the maximum job lateness (note that due dates are irrelevant for the makespan minimization). Suppose preemptions are allowed. Then we have a simple method to solve the problem optimally even on-line. We usually start at the earliest job release time and set initially the current time $t$ to that magnitude. Iteratively, among the available jobs, we schedule the next a most urgent job, one with the smallest due date. However, if during the execution of that job, another more urgent job gets released, the former job is interrupted at the release time of the latter job, which gets included at its release time. In this way, there will occur no forced delay for any urgent job by a non-urgent one, unlike the case without preemptions (in a given schedule $S$, the delay of job $j$ is $t_{j}^{S}-r_{j}$). The so constructed schedule is compact also in the sense that it minimizes the makespan. In particular, it has no avoidable machine idle time.
Assume that job $i$ is processed at time $r_{j}$ in the schedule $S$, where $d_{i}>d_{j}$. Without allowing the interruption of $i$, this job delays the starting of job $j$; note that this delay will be less than $p_{i}$ since upon the completion of job $i$, job $j$ can immediately be scheduled. The delay of job $j$ may result in a non-optimal schedule. Indeed, this will be the case when job $j$ is too urgent with a “small” due-date. In an optimal schedule, it should be included with no or with a smaller delay.
Now we describe a basic popular greedy non-preemptive algorithm to which we will refer as earliest due-date (ED) heuristic). It that has been widely used for solution of different single machine and other more complex scheduling problems. As the above described preemptive one, it is on-line. The heuristic was suggested by Jackson [31] and later developed by Schrage [45]. The first version is just the above described greedy algorithm with the pre-processing stage. The second version was adopted for jobs having both, release and due times. We describe this heuristic now. We let the current scheduling time $t$ be the minimal release time of an unscheduled job or the time when the machine completes the last job, whichever magnitude is larger. If no job is yet scheduled, we let $t:=0$. Iteratively, among the jobs already released by time $t$, the heuristic schedules one with the smallest due date. It breaks ties selecting a longest job.
Let $\sigma$ be the initial ED-schedule generated by the ED-heuristic for the originally given problem instance. Since no urgent job can be delayed by more than the processing time of another job, it is easy to see that
$L_{\max}(\sigma)<p_{\max}=\max_{j\in J}p_{j}.$
In particular, $\sigma$ is not necessarily optimal.
Let us introduce some additional concepts that will be useful to understand better the problem. Let $S$ be an ED-schedule, one generated by the ED-heuristic.
Among all jobs in the schedule $S$, it is useful to distinguish the ones whose lateness attains $L_{\max}(S)$. In a given block $B\in S$, let us call among them the last scheduled job from this block $B$ an overflow job in the schedule $S$. Let $o(S)$ be the earliest overflow job in the schedule $S$, and let the critical block $B(S)$ of $S$ be the block containing job $o(S)$.
Let now $e$ be a job from the block $B(S)$ such that $d_{e}>d_{o(S)}$, i.e., it is less urgent than the overflow job $o(S)$. Then job $e$ is referred to as an emerging job in the schedule $S$. The latest scheduled emerging job in the block $B(S)$ is called the delaying emerging job.
The kernel $K(S)$ in the schedule $S$ consists of the set of jobs, included in that schedule between the delaying emerging job $l(S)$ and the overflow job $o(S)$, not including job $l(S)$ (but including job $o(S)$). It immediately follows that the due date of any job of the kernel $K(S)$ is no less than that of job $o(S)$).
The above definitions can straightforwardly be applied to any (not necessarily critical) block. Then there may exist more than one overflow job (at most one in in each block) in a schedule. Note that we will have the same number of kernels in that schedule.
The earliest starting time of a kernel $K$ is given by
${\cal S}(K)=\min\{r_{j}|j\in K\}.$
The starting time ${\cal S}(K,S)$ of kernel $K$ in the schedule $S$ is the starting time of the first scheduled job of that kernel in the schedule $S$.
If the ED-schedule $S$ is not optimal, the overflow job $o(S)$ is to be restarted earlier. By modifying artificially the initial problem parameters and applying the ED-heuristic to this modified instance, an alternative ED-schedule in which job $o(S)$ is restarted earlier can be created, as we describe now. In the schedule $S$, we look for the overflow job $o(S)$, the kernel $K(S)$ and the set of all emerging jobs. In an attempt to reduce $L_{\max}(S)$ (restart job $o(S)$ earlier), we activate an emerging job $l$ for the kernel $K(S)$, that is, we reschedule job $l$ after the kernel $K(S)$. The activation is carried out in two steps. At step 1 the release time of job $l$ is increased artificially to it a magnitude, no-less than the largest job release time in the kernel $K(S)$. At step, ED-heuristic is applied to the modified in this way problem instance. Since now the release time of job $l$ is no less than that of any job of the kernel $K(S)$, and $d_{l}$ is larger than the due date of any job from the kernel $K(S)$, the ED-heuristic will give the priority to the jobs of the kernel $K(S)$ and will reschedule job $l$ after all former jobs. The ED-schedule $\{S\}_{l}$, obtained in this way, is called a (weakly) complementary to $S$ schedule.
Note that in the schedule $S$, some emerging jobs might have been originally included after the kernel $K(S)$. Then at least one such job will be included before the jobs of the kernel $K(S)$ in the schedule $\{S\}_{l}$ by the ED-heuristic (since no job of the kernel $K(S)$ is released at time $s_{l}(S)$). As a result, the kernel $K(S)$ may not be restarted at its earliest starting time. To guarantee the restarting of the kernel $K(S)$ at its earliest starting time, the above emerging jobs might be forced to remain after the kernel $K(S)$. To reach this goal, we increase the release times of these jobs, similarly as we did it for job $l$. The ED-schedule, obtained by the ED-heuristic for the modified in this way problem instance, is called a (strongly) complementary to $S$ schedule and is denoted by $S_{l}$. A complementary schedule can be abbreviated as a C-schedule.
Note that there is exactly one job less included before the kernel $K(S)$ in the schedule $S_{l}$ compared to schedule $S$. (The same condition does not held in $\{S\}_{l}$; in fact, we may have more jobs scheduled before the kernel $K(S)$ in the schedule $\{S\}_{l}$ than in the schedule $S$.) Then, as it is easy to see, the overflow job $o(S)$ will be completed earlier in the schedule $S_{l}$ than it was completed in the schedule $S$:
Lemma 2.1
Let $S$ be a strongly C-schedule and let $l$ be the emerging job scheduled in $i$th position in in that schedule. Let, further $j$ be the job, immediately succeeding job $l$ in schedule $S$. Then if $d_{j}<d_{l}$, there will occur a gap in between the $i$th scheduled job and the kernel $K(S)$ in the schedule $S_{l}$ Furthermore, there will arise a gap immediately before the earliest scheduled job of the kernel $K(S)$ in the complementary schedule $S_{l}$, where $l$ is the live emerging job in the schedule $S$. As a result, the kernel $K(S)$ will start at the minimum release time of a job from that kernel.
Proof. Assume first that job $j$ is scheduled in the $i$th position in the schedule $S_{l}$ and there arises no gap immediately before that job in that schedule. Then $r_{j}\leq r_{l}$ holds, but since $d_{j}<d_{l}$, the ED-heuristic would schedule job $j$ instead of job $l$ at the moment $t_{l}(S)$ in $S$, which is a contradiction. We use a similar reasoning if instead of job $j$ another job $k$ with $d_{k}<d_{l}$ is scheduled in the $i$th position in schedule $S_{l}$. If $d_{k}\geq d_{l}$, since the schedule $S_{l}$ is a strongly C-schedule, job $k$ must be an emerging job scheduled between job $l$ and kernel $K(S)$ in $S$. However, there is scheduled one job less before the kernel $K(S)$ in the schedule $S_{l}$ than in the schedule $S$, and our claim easily follows.
To distinguish the type of a gap in Lemma 2.1 from a “natural” gap that arises in the initial ED-schedule, the former type of a gap is sometimes referred to as an “artificial” one.
3 Polynomially Solvable Cases, Complexity and Approximability Issues
As earlier noted, the non-preemptive single machine scheduling problem becomes non-trivial, if, besides job the processing times, a job has at least one more parameter (as otherwise, from any permutation of the $n$ jobs, an optimal makespan schedule can be constructed in linear time by assigning the jobs from that permutation without leaving any idle time interval on the machine). In particular, all jobs are simultaneously released, any non-idle time schedule minimizes the makespan $C_{\max}$.
Minimizing maximum job lateness $L_{\max}$ is a bit more complex (recall that the lateness of job $j$ in a schedule $S$ is $L_{j}^{S}=c^{S}_{j}-d_{j}$). If only the job due dates are given (all the jobs being released at time 0), the Earliest Due date (ED) heuristic obtains an optimal schedule minimizing maximum job lateness (Jackson [31]). This heuristic schedules the jobs in non-decreasing order of their due dates without any idle time intervals (gaps). Vice versa, if only job release times are given, the Earliest Release time (ER) heuristic generates an optimal makespan schedule. In that heuristic, initially, the minimum job release time is assigned to the current scheduling time, and any job released at that time is assigned to the machine. Iteratively, the next scheduling time is defined as the maximum between the completion time of the latest so far assigned job and the minimum release time taken among the yet unscheduled jobs. Again, any (unscheduled) job released at the current scheduling time is assigned to the machine.
Because of the job release times, a schedule created by the ER-heuristic may contain a gap. However, it may contain no gap that might be avoided. In particular, the total gap length in such a schedule is the minimal possible one. We can also use this heuristic if all the jobs have a common due date $d$. In other words, all jobs are equally urgent. Then, as it is easy to see, an optimal schedule minimizing the maximum job lateness contains no gap that might be avoided, and it has the minimum possible total gap length. Therefore, the ER-heuristic minimizes also the maximum job lateness.
It is easy to see that the ER-heuristic will no more minimize the maximum job lateness even if only two possible job due dates are allowed. Indeed, consider an instance with two jobs, job 1 released at time 0 and having the due date 11 and job 2 released at time 1 with the due date 6. The ER-heuristic will assign job 1 to the time interval $[0,5)$ and job 2 to the time interval $[5,10)$ resulting in the maximum job lateness 4 (that of job 2). Observe that this schedule contains no gap. At the same time, an optimal schedule does contain a gap $[0,1)$: it assigns job 2 to the time interval $[1,6)$ and then job 1 to the time interval $[6,11)$ completing both jobs on-time with the maximum job lateness equal to 0.
3.1 Scheduling a Single Machine with Release Times and Due Dates to Minimize Maximum Job Lateness
The general problem with release tames and due dates to minimize maximum lateness, abbreviated as $1|r_{j}|L_{\max}$, is strongly NP-hard [24]. If we replace the due dates by tails with the objective to minimize maximum lateness, we get an equivalent problem $1|r_{j},q_{j}|C_{\max}$. In this transformation, every due date $d_{j}$ is replaced by the tail $q_{j}=A-d_{j}$, where $A$ is a sufficiently large constant (no less than the maximum job due date). So, the smaller is the due date, the larger is the tail. In other words, the urgency of every job is conserved with this transformation. It is straightforward to see that the two versions are equivalent. The full completion time of job $j$ in $S$ is given by ${\cal C}_{j}(S)=c_{j}^{S}+q_{j}$. The objective in that setting is to find an optimal makespan schedule, i.e., a feasible schedule $S$ with the minimal value of the maximal full job completion time.
An efficient implicit enumerative algorithms for the problems $1|r_{j}|L_{\max}$ and $1|r_{j},q_{j}|C_{\max}$ was suggested in McMahon and Florian [39] and later, based on similar ideas, in Carlier [13]. A number of other special cases are solvable in polynomial time. If all jobs are of unit length, the problem is again easy to solve, see Horn [30]: indeed, while applying the ED-heuristic, no new (earlier unreleased) job can be released within the execution interval of a currently running job since job release times are integer and every job takes just a unit time of processing. As a result, no job may cause a forced delay of any more urgent job.
As we have already seen, if all $r_{j}$s are equal, then scheduling the jobs in order of non-decreasing due dates gives an optimal schedule in $O(n\log n)$ time [31]. Similarly, if all $d_{j}$s are equal, then scheduling the jobs in order of non-decreasing release times is optimal. An $O(n\log n)$ solution of the problem was proposed in [28] for the special case when the release times, processing times and due-dates are restricted in a specific manner so that each $r_{j}$ lies within a certain time interval defined by $p_{j}$ and $d_{j}$.
With arbitrary release times and due dates (or tails), if all processing times are equal (but not necessarily of unit length), it is still possible to find an optimal schedule in polynomial time, see Garey et al. [25]. Even with two allowable job processing times $p$ and $2p$, for an arbitrary integer $p$, the problem $1|p_{j}\in\{p,2p\},r_{j},q_{j}|C_{\max}$ remains polynomially solvable [48]. In fact, in the algorithm from the latter reference the processing times of only some jobs (which become known during the execution of the algorithm) are restricted to $p$ or $2p$, whereas the rest of the jobs may have arbitrary processing times. The algorithm is obtained as a consequence of two enumerative algorithms for the problem $1|r_{j},q_{j}|C_{\max}$. The first algorithm is exponential. The the second algorithm is a restriction of the first one and runs in time $O(n^{2}\log n)$ under certain conditions which are to be held during its execution. Otherwise, another auxiliary procedure is to be applied to guarantee the optimality. An $O(n^{2}\log n\log P)$ implementation of that procedure for the problem $1|p_{j}\in\{P,2P\},r_{j},q_{j}|C_{\max}$ yields an algorithm with the same time complexity for this setting.
If the maximum job processing time $p_{\max}$ is bounded from above by a polynomial function in $n$ (i.e., $P(n)=O(n^{k})$) and the maximum difference between the job release times is bounded by a constant $R$, then the corresponding parameterized problem $1|p_{max}\leq P(n),|r_{j}-r_{i}|<R\ |L_{\max}$ can still be solved in polynomial time $O(n^{k+1}\log n\log p_{\max})$, see Vakhania & Werner [50]. Note that in this setting the general problem is parameterized by $p_{\max}$ and the maximum difference between two job release times.
There are known other conditions that allow a polynomial time solution of the problem. For example, if the job parameters are embedded in a special manner, i.e., for any pair of jobs $j,i$ with $r_{i}>r_{j}$ and $d_{i}<d_{j}$,
$d_{j}-r_{j}-p_{j}\leq d_{i}-r_{i}-p_{i}$
holds, and with
$r_{i}+p_{i}\geq r_{j}+p_{j},$
$d_{i}\geq d_{j}$ holds, then the resultant problem $1|r_{j},embedded|L_{\max}$ is solvable in $O(n\log n)$ time, see Vakhania [52].
In a maximal polynomially solvable special case of the problem $1|p_{j}:divisible,r_{j}|L_{\max}$, the job processing times form mutually divisible numbers (say, powers of 2) [53].
For the problem $1|r_{j},p_{j}=p|L_{max}$ with constant processing times, Lazarev et al. [38] presented two polynomial algorithms. The first procedure is based on trisection search and determines an optimal solution in $O(Qn\log n)$ time, where $10^{-Q}$ represents the accuracy of the parameters of the problem. The second approach uses an auxiliary problem, where the maximum completion time is minimized under the constraint that $L_{max}$ is bounded by a particular value. Using this approach, the authors determine for the bi-criteria problem with the two criteria $C_{max}$ and $L_{max}$ and the Pareto set of optimal solutions in $O(n^{3}\log n)$ time.
With arbitrary job processing times, the problem remains (weakly) $NP$-hard already with only two allowable job release times and due dates $1|r_{j}\in\{r_{1},r_{2}\},d_{j}\in\{d_{1},d_{2}\}|L_{\max}$, see Chinos and Vakhania [14]. This is a minimal $NP$-hard special case of the problem $1|r_{j},d_{j}|L_{\max}$. The latter setting was later studied in Reynoso and Vakhania [44], where structural properties of this special case have led to a number of optimality conditions under which it can be solved optimally in $O(n\log n)$ time were established. It was shown that, for bottleneck instances where none of these conditions apply, the problem can be reduced to the SUBSET SUM problem. The authors presented computational results accomplished for 50 million randomly generated problem instances, reporting that almost all these instances were solved optimally by one of the established optimality conditions. Moreover, the heuristic algorithms that incorporate the optimality conditions established in [44] succeeded to solve the SUBSET SUM problem also optimally for almost all the 50 million instances. A very recent work considered a natural generalization of the latter problem with any fixed number of release times, see Escalona and Vakhania [17]. For this setting, the optimality conditions were derived and again, it was shown how the problem can be reduced to the SUBSET SUM. It is a challenging question to what extent the above obtained results can be generalized for any fixed number of release and delivery times, a setting with much wider practical applications.
In [7] a more general objective function $f_{max}$ defined as the maximum of non-decreasing cost functions $f_{j}$ has been considered for the preemptive problem $1|r_{j},prec,pmtn|f_{max}$. It is assumed that the cost functions fulfill the order relation $f_{i}(t)\leq f_{j}(t)$ or $f_{j}(t)\leq f_{i}(t)$ for any pair $i,j$ of jobs and for any time $t$ of the scheduling interval. In the case of tree-like precedence constraints, the derived algorithm has a time complexity of $O(n\log n)$. In addition, a parallel algorithm is given with a time complexity of $O(\log^{2}n)$ when using $O(n^{2}\log n)$ processors.
A special kind of parameterized analysis called the Variable Parameter (VP) analysis was recently suggested in [54]. Based on this approach, an implicit enumeration algorithm and a polynomial time approximation scheme for the general setting $1|r_{j},q_{j}|C_{\max}$ were proposed. The variable parameter is defined as the number of emerging jobs. First, a partial solution in time $O(n\log n)$ without these jobs is constructed. Then this partial solution is augmented to a complete optimal solution in time, which is exponential solely in the number of the variable parameter (the number of emerging jobs). The author also gives alternative time complexity expression for both, the implicit enumeration and the polynomial time approximation scheme, in which the exponential dependence is solely on the maximum processing time and the maximum delivery time of an emerging job. If one applies the fixed parameter analysis to these estimations, one gets a polynomial-time dependence. A very recent extensive experimental study showed that, in practice, the variable parameter is much less than $n$. Moreover, the ratio of the variable parameter to $n$ asymptotically converges to 0 (see Ramirez and Vakhania [43]).
The problem $1|r_{j}|L_{\max}$ is important on its own and also because of its possible applications for the solution of other more complex scheduling problems and other optimization problems. For example, in the Multiple Traveling Salesman Problem (M-TSP), $k$ salesmen, each starting from a special point 00 called depot, has to visit a number of clients and return to the depot. The objective is to minimize the total cost of all $k$ tours (see, for example [42]). In the setting with time windows M-TSP-TW, each client has to be visited within a specified time interval. In the former basic setting, it is easy to find a feasible solution to the problem. However, in the latter setting with time windows, there may exist no feasible solution for a given $k$. Hence, it is crucial to determine whether there exists a feasible solution for a given $k$. More generally, one may be interested to calculate the minimum $k$ for which there exists a feasible solution. We can use our scheduling problem for this purpose.
In the basic case, we wish to determine whether there exists a feasible solution with $k=1$. For a given instance of M-TSP-TW with $n+1$ clients, we create an instance of he problem $1|r_{j},p_{j}=0|L_{\max}$ with $n$ jobs, associating job $j$ with client $j$, for each $i=1,\dots,n$, and letting the processing time of each job to be 0. Note that this is a simplified version of the generic problem $1|r_{j}|L_{\max}$. In this version, the release and due times of each job are determined according to the time window of the corresponding client. We consider the feasibility version of the problem, where one looks for a schedule in which no job has a positive lateness. One can immediately see that, if there is no feasible solution to the scheduling problem, there is also no feasible solution for the corresponding instance of the problem M-TSP-TW with $k=1$ (i.e., TSP-TW). In general, we let $u$ be an upper bound on the total number of salesmen, and carry out binary search within the interval $[1,u]$ solving at each iteration an instance of the corresponding scheduling problem $Pk|r_{j},p_{j}=0|L_{\max}$ with $k$ parallel identical machines and with trial $k\in[1,u]$. In this way, we can find the minimum $k$ for which there exists a feasible solution to the problem M-TSP-TW.
3.2 Minimizing the Number of Late Jobs
Another common objective function with given due dates is the number of late jobs: a job in a schedule is called late if it is completed after its due date in that schedule, otherwise it is called on time. In the problem of this subsection, we deal with $n$ jobs with release times and due dates which have to be scheduled on the machine so as to maximize the number of on time jobs, or equivalently, to minimize the number of late jobs. We consider both, preemptive and non-preemptive cases, and also cases when the jobs have an additional parameter called weight.
3.2.1 Nonpreemptive Scheduling to Minimize the Number of Late Jobs: $1|r_{j};p_{j}=p|\sum U_{j}$
In this subsection, we deal with the non-preemptive version of the problem $1|r_{j}|\sum U_{j}$. This problem was studied in Vakhania [51], where two polynomial time algorithms were proposed. The first one with the time complexity $O(n^{3}\log n)$ solves optimally the problem if during its execution no job with some specially defined property occurs. The second algorithm is an adaptation of the first one for the special case of the problem when all jobs have the same length $1|r_{j};p_{j}=p|\sum U_{j}$. The time complexity of this algorithm is $O(n^{3}\log n)$ (it improved an earlier known fastest algorithm with the time complexity of $O(n^{5})$ by Chrobak et al. [15].
3.2.2 Preemptive Scheduling to Maximize the Number of on Time Jobs: $1|r_{j};pmtn|\sum U_{j}$
In the problem of this subsection, $n$ jobs with release times and due dates have to be scheduled preemptively on a single machine so as to minimize the number of late jobs, i.e., problem $1|r_{j};pmtn|\sum U_{j}$. We call a job late (on time, respectively) if it is completed after (at or before, respectively) its due date. We will say that job $j$ is feasibly scheduled if it starts no earlier than at time $r_{j}$ and overlaps with no other job on the machine. A feasible schedule is one which feasibly schedules all jobs. The objective is to find a feasible schedule with a minimal number of late jobs, or equivalently, with the maximal number of on time jobs.
This problem is known to be polynomially solvable. The first polynomial dynamic programming algorithm with a time complexity of $O(n^{5})$ and a space complexity of $O(n^{3})$ was developed by Lawler [33]. The result was improved by Baptiste [8] who suggested another dynamic programming algorithm with a time complexity of $O(n^{4})$ and a space complexity of $O(n^{2})$. The latter result was further improved by Vakhania [49], where an algorithm with the time and space complexities of $O(n^{3})$ and $O(n)$, respectively, was proposed. Unlike the earlier algorithms, the latter algorithm is not based on dynamic programming. It uses the ED-heuristic for scheduling the jobs. The main task is to construct a schedule in which all jobs are scheduled on time, with the number of such jobs being maximal. As any late job can be scheduled arbitrarily late without affecting the objective function, jobs not included in the former schedule can be later added at the end of it in an arbitrary order. During the construction of the schedule, some jobs are finally disregarded, whereas some other jobs are omitted with the possibility to be included into a later created schedule.
3.3 Preemptive Scheduling to Minimize the Weighted Number of Late Jobs with Equal Length: $1|r_{j};p_{j}=p;pmtn|\sum w_{j}U_{j}$
In this section, we consider the weighted version of the problem of the previous section in which preemptions are allowed and all jobs have an equal length. Moreover, $w_{j}$ is the non-negative integer weight of job $j$. We minimize the weighted throughput which is the total weight of the completed jobs. Equivalently, we minimize the weighted number of late jobs. The problem ican be abbreviated as $1|r_{j};p_{j}=p;pmtn|\sum w_{j}U_{j}$.
The first polynomial dynamic programming algorithm of the complexity $O(n^{10})$ for the problem was suggested by Baptiste [9]. Baptiste et al. [11] improved the above time complexity and proposed another dynamic programming algorithm which runs in $O(n^{4})$ time.
Next, we briefly discuss some directions for further research. First we note that for the multi-processor case, the weighted version is $NP$-complete [12]. At the same time, the non-weighted version still remains open. In particular, it is not known whether the problem $P|r_{j};p_{j}=p;{pmtn}|\sum U_{j}$ can be solved in polynomial time. (Even for $2$ processors, one cannot the search to solely earliest deadline schedules. As an example, consider an instance with three jobs with feasible intervals $(0,3)$, $(0,4)$, and $(0,5)$ and the processing time $p=3$. There is a feasible schedule that completes all three jobs, whereas the earliest-deadline schedule completes only jobs $1$ and $2$.) For the multi-processor case, it is also interesting to study the preemptive version, where jobs are not allowed to migrate between processors.
3.4 Scheduling to Minimize the Weighted Number of Late jobs with Deadlines and Release/Due Date Intervals
In [26], Gordon et al. dealt with the problem, where in addition to release and due dates also deadlines are given. This means that certain jobs have to be completed on time. The authors considered the two cases of similarly ordered release and dues dates (i.e., $r_{1}\leq r_{2}\leq\ldots\leq r_{n}$ and $d_{1}\leq d_{2}\leq\ldots\leq d_{n}$ as well nested release and due date intervals (i.e., none of two such intervals do overlap, this they have at most joint point or one covers the other one). For these problems, a reduced problem is constructed, where the deadlines are eliminated, and it is shown that the optimal solution of the reduced problem can be used to find an optimal solution of the initial problem.
In a later paper, Gordon et al. [27] considered the preemptive version of the above problem with nested release and due date intervals. They derived necessary and sufficient conditions for the existence of a schedule in which all jobs are completed by their due dates. For the case of oppositely ordered processing times and job weights, a polynomial algorithm of the complexity $O(n\log n)$ has been presented.
3.5 Scheduling with Tardiness-based Objectives
While the problem $1||\sum T_{j}$ is known to be NP-hard, Gafarov et al. [22] considered the special case of equal processing times, i.e., problem $1|r_{j},p_{j}=p|\sum w_{j}T_{j}$. After deriving some properties, they gave polynomial algorithms for two special cases of the problem. If the difference between the maximum due date and the minimum due date is not greater than the processing time $p$, the problem can be solved by a modification of the algorithm by Baptiste [10] in $O(n^{10})$ time. The special case with ordered weights and processing times according to $w_{1}\leq w_{2}\leq\ldots\leq w_{n}$ and $d_{1}\leq d_{2}\leq\ldots\leq d_{n}$ can be solved in $O(n^{9})$ time by a similar modification from [10].
In [36], Lazarev and Werner presented a new general graphical algorithm which often improves the complexity or at least the running time of dynamic programming algorithms, and it can be used for efficient approximation algorithm. Compared to standard dynamic programming algorithms, the number of states can often substantially be reduced.
In a standard dynamic programming algorithm for permutation problems, functions $F_{l}(t)$ and the corresponding partial sequences have to be stored for any stage $l=1,2,\ldots,n$ and any time $t=0,1,\ldots,UB$. Here $UB$ is an upper bound for the scheduling interval, e.g. the sum of all processing times (or the maximal due date). The values $F_{1}(t)$ typically represent the best function value when considering the first $l$ jobs and allowing at most $t$ time units for the processing of the scheduled (e.g. early) jobs. This results in a pseudo-polynomial algorithm of the complexity $O(n\sum p_{j})$
The graphical algorithm combines different states, where the best partial solution does not change and the corresponding objective function $F_{l}(t)$ can be described by the same linear function. As a result, at each stage $l$, there are certain intervals separated by the so-called breakpoints $t^{l}_{1}<t^{l}_{2}<\ldots t^{l}_{h_{l}}$, where the structure of the linear function belonging to the best partial solution changes. In addition, for the graphical algorithm, the times $t$ do not need to be integer.
In [23], Gafarov et al. described the application of this graphical algorithm to the approximate solution of various single machine problems. In particular, the graphical algorithm is applied to 5 total tardiness problems, where a pseudo-polynomial algorithm exists, and the graphical algorithm yields fully polynomial-time approximation schemes yielding the best known running time.
In [21], Gafarov et al. considered the maximization of total tardiness and by applying the graphical algorithm, they transformed a pseudo-polynomial algorithm into a polynomial one. More precisely, the complexity of $O(n\sum p_{j})$ can be reduced to $O(n^{2})$ what settled the open complexity status of the problem $\max 1|(nd)|\sum T_{j}$), where $nd$ means that no inserted idle times occur. The authors present a detailed numerical example and give also a practical application of such maximization problems, where wind turbines have to be mounted in several regions of a country.
While for more general settings, such a graphical algorithm improves often the running time in comparison with existing algorithms, the complexity remains pseudo-polynomial. For instance, in [20], Lazarev et al. presented graphical algorithms with a complexity of $O(n\sum p_{j})$ for a special case of a generalized tardiness problem corresponding to the minimization of late work as well as for the minimization of the weighted number of late jobs.
In [37], Lazarev and Werner considered various special cases of the (unweighted) total tardiness problem $1||\sum T_{j}$. After proving that already the special case of oppositely ordered processing times and due dates, i.e. $p_{1}\geq p_{2}\geq\ldots\geq p_{n}$ and $d_{1}\leq d_{2}\leq\ldots\leq d_{n}$, is $NP$-hard in the ordinary sense, they gave several pseudo-polynomial and polynomial algorithms for special cases of this problem. In particular, $O(n^{2})$ algorithms have been given for the cases $d_{j}-d_{j-1}>p_{j},j=2,3,\ldots,n$ as well as for the case that the difference between the maximal and the minimal due date is not greater than 1. In both cases, the processing times do not need to be integer.
In further research, the graphical algorithm can be applied to more general problems, where a pseudo-polynomial algorithm exists and Boolean variables are used for making yes/no decisions.
3.6 Preemptive Scheduling in Overloaded Systems
In the single machine scheduling problem in overloaded systems, each job has a release time, a deadline and a weight. The latter parameter reflects its profit. Any job can be preempted during the execution, and the objective is to maximize the overall profit. Two possibly models and the definition of the profit gained by the execution of task $j$ can be defined. In the standard model, a completed task $j$ gives the profit $w_{j}p_{j}$, whereas a non-completed task gives no profit. In the metered model, a task $j$ executed for $t\leq p_{j}$ time units gives the profit $w_{jt}$ (this task may not be completed). For the off-line metered profit model, Chrobak et al. [16] used a reduction to bipartite matchings and maximal flows to develop a polynomial-time algorithm. For the online metered case, they presented an algorithm with competitive ratio $e/(e-1)\approx 1.5820$ and proved a lower lower bound of $\sqrt{5}-1\approx 1.236$ on the competitive ratio of algorithms for this problem.
3.7 Scheduling with Restarts
In the on-line single machine scheduling problem with restarts, jobs arriving over time are characterized by processing, release and delivery times, which become known at the arrival (release) time of each job. The objective is to minimize the maximum full job completion time, the makespan. As earlier mentioned, the corresponding off-line problem is strongly NP-hard, but the preemptive version can be solved in polynomial time by the preemptive LDT-heuristic. For the non-preemptive case, the LDT-heuristic has a performance guarantee of 2, which is the best possible one unless we allow the machine to be idle when there is an unprocessed job available. Hoogeveen and Vestjens [29] developed an on-line algorithm with the best possible worst-case ratio
$(\sqrt{5}+1)/2\approx 1.61803$
in which a waiting strategy is incorporated. Van den Akker, Hoogeveen and Vakhania [3] considered an extension of the standard model by allowing job restarts: if an urgent job comes in while the machine is busy, then the currently running job can be interrupted in favor of the former urgent job. In this model, unlike traditional preemptive models, the already processed part of the interrupted job is wasted, this job is to be started all over again. The algorithm proposed in [3] has the best possible worst-case ratio of $1.5$ (which is better than the above mentioned bound for the case without the restarts).
Again, we give some possible directions for further research in connection with allowed restarts. As we just mentioned, with the restart model considered in [3], if a job is restarted, then the portion of this job that was already processed by the machine is lost. This may be avoided for some applications, where the already completed part of an abandoned job gives some benefit and this job can be restarted from the time moment when it was interrupted or at some earlier time. Hence, one may consider models with partial restarts when the already processed part of an interrupted job counts partly or completely. In particular, it is interesting to find out whether there exist better worst-case performance bounds for models with partial restarts.
3.8 Scheduling with Financial Resource Constraints
Gafarov et al. [19] considered single machine problems with a non-renewable resource (abbreviated $NR$) like e.g. money or fuel. This means that at any time $t_{i},i=0,1,\ldots,UB$ with $t_{0}=0$ and $UB$ as an upper bound on the scheduling interval, one gets an amount $G(t_{i})\geq 0$ of the resource, while any job $j$ has a consumption of $g_{j}$ at the starting time $s_{j}$ of the job. Hence, the equality
$\sum_{j=1}^{n}g_{j}=\sum_{i=0}^{UB}G(t_{i})$
holds. One looks for an optimal permutation $(j_{1},j_{2},\ldots,j_{n})$ for a particular objective function with the starting times $s_{j_{1}},s_{j_{2}},\ldots,s_{j_{n}}$ of the jobs such that the resource constraints are satisfied, i.e., the inequality
$\sum_{l=1}^{i}g_{j_{l}}\leq\sum\{G(t_{v})\mid t_{v}\leq s_{j_{i}}\mbox{ for all }v\leq s_{j_{i}}\}$
holds for any $i=1,2,\ldots,n$. Such problems play a role e.g. for government-financed organizations when a project can only be started after receiving the required funds.
The authors focused on complexity results and proved that the problems $1|NR|f$ with $f\in\{C_{max},\sum C_{j},\sum U_{j}\}$ are strongly $NP$-hard. When minimizing total tardiness $\sum T_{j}$, the case of equal due dates ($d_{j}=d$) is already strongly $NP$-hard. The following special cases with minimizing total tardiness were proven to be weakly $NP$-hard:
1) equal processing times ($p_{j}=p$);
2) equal due dates $(d_{j}=d)$ and equal consumptions ($g_{j}=g$).
On the other hand, the special case of equal processing times ($p_{j}=p$) and ordered consumptions and due dates according to $g_{1}\leq g_{2}\leq\ldots\leq g_{n}$ and $d_{1}\leq d_{2}\leq\ldots\leq d_{n}$ is polynomially solvable.
This paper considered also the budget scheduling problem with the objective to minimize the makespan. Instead of a resource consumption $g_{j}$ for each job, two values $g_{j}^{-}$ and $g_{j}^{+}$ are given, where $g_{j}^{-}$ has the same meaning as $g_{j}$ before. The value $g_{j}^{+}$ describes the additional earnings at the completion time of job $j$. The authors showed that the problem $1|NR,g_{j}^{-},g_{j}^{+}|C_{max}$ belongs to the class $APX$ of problems allowing a constant-factor approximation algorithm.
For future work, due to the $NP$-hardness of the majority of problems, it would be interesting to derive elimination rules and structural properties of optimal solutions which can be used in enumerative algorithms.
4 Concluding Remarks
In this paper, we reviewed some theoretical results on classical single machine scheduling problems. Since there exists a huge number of heuristic/metaheuristic and enumerative algorithms, our focus was on polynomial algorithms as well as complexity and approximability issues, discussed in the papers authored or co-authored by the authors of this work. The summarized polynomial special cases are also of interest when the conditions are not exactly satisfied and the resulting solution can be taken as a good heuristic one.
For single machine scheduling problems, efficient solution methods can often be found easier than for the corresponding multiprocessor or shop scheduling problems. However, such methods may give a better insight into the latter more complex problems and may well serve as a basis for the development of solution methods for these problems. At the same time, optimal solutions to the single machine scheduling problems can directly be used to solve the corresponding multiprocessor or shop scheduling problems. For example, for such $NP$-hard problems, strong lower bounds can be obtained based on the solutions of the corresponding single machine problems. Non-strict lower bounds, which are easier and faster to obtain, can be used in any search-tree based approximation algorithm, such as beam search, for example. Similarly, using strict lower bounds, we obtain exact branch-and-bound algorithms. Once implemented and tested, the above algorithms can be used as an algorithmic engine for the construction of decision support systems. Finally, we note that a number of algorithms for basic single machine problems can relatively easily be adopted using additional graph-completion mechanisms to solve the corresponding extended versions with precedence relations, transportation and setup times and other real-life job scheduling problems.